Ramesh crosses a street 600 m long in 5 minutes. His speed in km/
hr is

Answer by mananth(16946)

Ramesh crosses a street 600 m long in 5 minutes. His speed in km/
hr is
600 m is = 0.6 km.
5 minutes = 5/60 hrs= 0.08 hours
speed = distance / time
= 0.6 / 0.08
his speed is 7.5 kms/ hr